OpenStudy (issy14):

Derive: 30. f(s) = 5^s * e^s Very confused here, wouldn't e^2 be a constant? shouldn't that be 0? then it would just be ln(5) * 5^s? where or why am I wrong?

OpenStudy (ikram002p):

is it e^2 or e^s ?

OpenStudy (issy14):

e^s

OpenStudy (ikram002p):

Hint :- \(\large (f(x)g(x))'=f(x)'g(x)+f(x)g'(x)\)

ganeshie8 (ganeshie8):

\[\large 5^s = e^{s\ln 5}\]

OpenStudy (issy14):

we haven't learned that yet, we can't use the product or quotient rule

OpenStudy (ikram002p):

ok then use what ganesh said :)

OpenStudy (issy14):

Can you describe what you did please

OpenStudy (issy14):

I have no idea why you combined them or the reasoning behind the answer

ganeshie8 (ganeshie8):

cos differentiating a single exponent function is easier than differentiating the product : \(\large (e^{ax})' = a e^{ax}\)

ganeshie8 (ganeshie8):

I see that you didn't get why \(\large 5^s\) equals \(\large e^{s\ln 5}\)

OpenStudy (issy14):

exactly, I don't know why that's equivalent, I get that 5^s*e^s = (5e)^s then it would be the ln(5e) * 5e^s

ganeshie8 (ganeshie8):

that looks perfect !!

OpenStudy (issy14):

is that correct? then what would the next step be? that's where I'm stuck, can I factor it ?

ganeshie8 (ganeshie8):

you have used \(\large (a^x)' = a^x \ln a\) formula right ?

OpenStudy (issy14):

yes

ganeshie8 (ganeshie8):

\[\large (5^s * e^s)' = \left((5e)^s \right)' = (5e)^s \ln (5e)\] there is no next step, you're done !

OpenStudy (issy14):

really? but (BOB) back of the book says something else

ganeshie8 (ganeshie8):

what does it say

OpenStudy (issy14):

it has (1+ln(5)) 5^s e^s?

ganeshie8 (ganeshie8):

haha the book simplified \( \ln(5e)\)

ganeshie8 (ganeshie8):

\(\ln (ab) = \ln a + \ln b\)

OpenStudy (issy14):

aaaaaaaaaaaaaaaaaaaaaaaaaaaaaaa!!!!!!!!!!!! jesus christ, I'm going to have to borrow your brain for my test

ganeshie8 (ganeshie8):

lol, in test we can leave the answer as \((5e)^s \ln (5e)\) you won't loose marks for not simpilifies this

ganeshie8 (ganeshie8):

ln(5e) looks simple to me than (1 + ln5) one term is better than two terms :P

OpenStudy (issy14):

BOB and I don't get along very well sometimes. thank you very much. =).

ganeshie8 (ganeshie8):

yw :) good luck wid the test !

OpenStudy (issy14):

thanks =)